OpenStudy (anonymous):

algebraically find, in simplest form, half the exact area of an equilateral triangle whose perimeter is 36√5 help me please ? :D

OpenStudy (anonymous):

do u have any ideas?

OpenStudy (anonymous):

equilateral triangle has three equal sides divide 36√5 by 3 = 12√5 so each side is 12√5 

OpenStudy (anonymous):

yea i got that far but dont i need the height to find area? how can i get that?

jigglypuff314 (jigglypuff314):



OpenStudy (anonymous):

oh wow i didnt even see that

OpenStudy (anonymous):

so i would set 6 radical 5 equal to n then do 6 radical 5 times radical 3?

jigglypuff314 (jigglypuff314):

yes :) then (6√5)(√3) = 6√(5*3) = 6√15

OpenStudy (anonymous):

then just plug into area formula? :D

jigglypuff314 (jigglypuff314):

yep :) and remember that you're finding "half the exact area of an equilateral triangle"

OpenStudy (anonymous):

how owuld 12 radical 5 times 6 radical 15 work out? im not to good at multiplying radicals :(

OpenStudy (anonymous):

i dont think im doing it correctly

OpenStudy (anonymous):

72 radical 75?

jigglypuff314 (jigglypuff314):

like (√5)*(√15) = √(5*5*3) = 5√3

jigglypuff314 (jigglypuff314):

yes, that too :)

OpenStudy (anonymous):

i would have to simplify the radical 75 though?

jigglypuff314 (jigglypuff314):

mmm I think it would be easier if you do

OpenStudy (anonymous):

216 radical 5?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

360 radical 3?

jigglypuff314 (jigglypuff314):

yeah and remember triangle formula had a (1/2) somewhere :P

OpenStudy (anonymous):

yes so 180 radical 5 then divide hat by 2 again and get 90 radical 5 as answer?

OpenStudy (anonymous):

radical 3

jigglypuff314 (jigglypuff314):

yes, I think so :)

OpenStudy (anonymous):

ok thank u so much :D

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

ima become your fan :D

jigglypuff314 (jigglypuff314):

Thank you :)

OpenStudy (anonymous):

thanks again i have to go , have a good night